# Remote Hawking-Moss instanton and the Lorentzian path integral

### v1

@@ I Introduction
Vacuum phase transition is one of the most important phenomena in cosmology. The decay of a false vacuum is a quantum-tunneling process of a field $\phi$ from a metastable false vacuum to a stable true vacuum in the effective potential $V(\phi)$ at lower or zero temperatures. This type of phase transition was first discussed in Refs. [1, 2], where the transition rate was evaluated by computing the on-shell Euclidean action, which corresponds to the non-trivial configuration of the field in the Wick-rotated space. Most of the subsequent studies have been conventionally following the Euclidean path integral proposed in Refs. [1, 2]. One of the simplest models of the transition in curved spacetime derived using the Euclidean path integral method is the Coleman-De Luccia (CDL) bounce [3], which respects the $O(4)$ symmetry. When a positive large cosmological constant governs the background and a potential barrier is shallower, stochastic fluctuations can be dominant in the transition process. Indeed, in de Sitter (dS) background, the quantum fluctuations of long wavelength modes in a dS patch are squeezed and may be treated as stochastic fluctuations [4, 5, 6, 7], which is often referred to as “the quantum-to-classical transition” of fluctuation. The most famous instanton describing such a transition is known as the Hawking-Moss (HM) instanton [8, 9]. The transition amplitude of the HM bounce can be indeed reproduced by solving the Fokker-Planck equation in an equilibrium condition [10]. Such a process is often interpreted as a thermal hopping process [11, 12, 13, 14].
@@
In Ref. [11], it was argued that the HM amplitude can be understood as the Boltzmann factor, i.e., a probability of finding the homogeneous configuration at the top of maxima in the associated temperature [11, 12, 13]. This interpretation should be valid only after the whole configuration is in a thermal equilibrium state. For the remote HM transition, it is not clear which thermal or quantum-tunneling effect is more important. As a non-equilibrium state, which could be involved in the case (I) as we will discuss later, is hard to describe in general, we here consider the contribution of the quantum tunneling effect to the remote HM transition. Also, the thermal effect would be suppressed when we address the issue (II). We will then study issues (I) and (II) by considering only the quantum tunneling effect while turning off the thermal/stochastic noises.
In this paper, we discuss the two issues raised above by utilizing the Lorentzian path integral with the Picard-Lefschetz theory. Recently, a method to evaluate the transition amplitude in the Lorentzian path integral was proposed in the context of quantum cosmology [18]. In Ref. [18], the authors applied the Picard-Lefschetz theory to find the relevant saddle point in the action, which is important to make the integration convergent. This important technique to perform the Lorentzian path integral was also applied to tunneling processes such as the radio astronomy [19], the tunneling problem in quantum mechanics [20, 21], the thin-shell vacuum decay process [22] and so on ([23, 24, 25]). We use this technique to see the quantum tunneling effect in the up-tunneling process. One may think that the quantum tunneling effect on the homogeneous transition is negligible as the primordial fluctuation may be decohered on the superhorizon scale. However, the decoherence effect may highly depend on how we split the whole system into an observed system and environment, which involves some artificial choice, more or less. In this sense, the quantum nature of the fluctuation may not vanish even in the superhorizon, and it could affect the subhorizon dynamics. The detailed discussion on this is beyond our scope in this paper.
@@ II Euclidean formalism: CDL bounce and HM bounce
where we have used Eq. (14). Eq. (18) implies that the transition amplitude of the HM bounce is governed by the Boltzmann factor associated with the Bekenstein-Hawking temperature of the dS horizon. From this perspective, the transition amplitude for the HM bounce can be understood as the probability of finding the homogeneous configuration of the scalar field at the maxima of the potential in the thermal system of $T=T_{H}$. In that sense, the HM bounce is often interpreted as a “thermal” transition. Also, as this Boltzmann factor was obtained via the Fokker-Planck equation with the stochastic noise of a homogeneous field in the dS spacetime [10], we assume that the stochastic noise can be identified as the thermal fluctuations. On the other hand, we will here consider the quantum tunneling effect, which is distinct from the thermal effect, to get some insight into the remote HM bounce Ref. [11, 12].
@@ IV.2 Classical action and saddle point
Before presenting our results, it should be noted how we model the remote HM transition in our setup. According to Ref. [11], the HM amplitude can be interpreted as a transition amplitude when we consider the transition from a false vacuum to the neighboring potential top and the potential is sufficiently flat near the top ($\gamma\ll 1$). Additionally, Ref. [11] assumes that the system is in thermal equilibrium. To ensure the system is thermalized within the relevant time scale, say $t$, the field distance between the false vacuum and the potential top $\Delta\phi$ should not be so long that the field cannot reach the top from the false vacuum within $t$. Assuming that the typical relaxation time to achieve equilibrium state, $t_{\rm eq}$, can be estimated by the time scale of the stochastic dissipation process of the field, one could use the following relation [27, 28, 29]
@@
where $H$ is the Hubble parameter and $\sim Ht_{\rm eq}$ is the step number of the stochastic process. One can read $t_{\rm eq}\sim(\Delta\phi)^{2}/H^{3}$ from (38) and find $t_{\rm eq}\propto 1/\gamma$ provided that $\Delta V$ and $H$ are fixed. As the conventional HM transition cannot be applicable to a situation with $t\ll t_{\rm eq}(\Delta\phi,H)$, there should be a certain threshold value $\gamma=\gamma_{\textmd{th}}\ll 1$ above which (i.e., $\gamma_{\textmd{th}}<\gamma\ll 1$) the field space around the false vacuum and potential top achieves equilibrium state within the relevant time scale. On the other hand, with a shallower potential ($\gamma<\gamma_{\textmd{th}}$) and a large $\Delta\phi$, thermal equilibrium is unlikely to be realized around the top of the barrier within the relevant time scale. Thus, if we want to discuss the issue (I) concerning the quantum transition to the remote potential top, the relevant situation is $\gamma<\gamma_{\textmd{th}}$ with $\Delta V$ being fixed. On the other hand, for issue (II) concerning quantum tunneling with a steeper barrier ($|V^{\prime\prime}|\gtrsim H^{2}$), it is reasonable to consider the region $\gamma\gtrsim 1$, assuming the stochastic noise can be neglected compared to quantum diffusion. In the following, based on the above discussion, we will evaluate the transition amplitudes for the two situations (I) and (II).
@@
Figure 3: The contour of $\text{Re}[iS_{0}(N)]$ for $\Delta V/V_{0}=10^{-6}$. $\gamma=10$ (upper left), $1$ (lower left), and $10^{-2}$ (lower right), respectively. In the upper left figure, the pink and black markers denote the relevant and irrelevant saddle points, and green/pink curves represent the steepest ascent/decent paths, respectively. The upper right figure shows $\text{Im}[iS(N)]$ for $\gamma=10$. In the upper left panel, the saddle points are depicted as a black diamond and pink circle. The pink solid arrow and green dashed arrow mean the steepest descent and ascent path from the lower saddle point (pink circle), respectively.
@@
We present the contour plots of $iS_{0}$ in Fig. 3. We can see that saddle points exist in the complex-$N$ plane. We find two saddle points close to the original contour at $(0,\infty)$ on the real-$N$ axis. Those saddle points are marked in the left top panel of Fig. 3. In the following, we shall denote the saddle points in the upper and lower half region in the complex-$N$ plane as the upper and lower saddle points, respectively. We observe that the steepest descent curves from the upper saddle point, drawn as an upward green dashed arrow, coincide with the steepest ascent curves from the lower one. This prevents us from making the one-to-one correspondence between the saddle point and the associated steepest descent and ascent contours. To resolve this degeneracy, an infinitesimal perturbation is introduced, as discussed in Refs. [18, 30]. Henceforth, we shall assume the perturbation has been added and disregard the degeneracy. Considering the deformation of the original contour $(0,\infty)$, we observe that only the lower saddle point contributes to the dominant transition.
@@ IV.4 Transition rate
In both cases, $\Delta V/V_{0}=10^{-6}$ and $10^{-3}$, shown in Fig. 5, the exponent $\text{Im}(S_{\rm eff})$ is equal to or higher than that of the HM bounce for $\gamma<\gamma_{\textmd{cr}}$. As $\gamma\ll 1$ corresponds to the remote HM transition, it is subtle if the conventional HM amplitude $e^{-{\cal B}_{\rm HM}/\hbar}$ gives the transition amplitude in that limit as discussed in Ref. [11]. According to Ref. [11], for the remote HM transition, the conventional HM amplitude merely means the probability of finding the homogeneous configuration, but this says nothing about the relative likelihood with the initial configuration. Thus, when we focus on the homogeneous transition from the false vacuum to the potential top, especially the distant one, the rate may be determined by the quantum tunneling amplitude rather than the Boltzmann factor as the dominant contribution. If this is the case, the remote HM transition would be suppressed as one can see in Fig. 5.
@@ V Conclusions and discussions
In this paper, we investigated homogeneous tunneling using the Lorentzian path integral, following the Picard-Lefschetz theory, and evaluated the tunneling probability of a homogeneous scalar field. We then discussed the following two issues: (I) the contribution of the quantum tunneling effect to the remote HM transition and (II) the HM transition with a steeper barrier for which the thermal effect would be suppressed. To this end, we turned off thermal/stochastic noises and took into account only the quantum tunneling effect by using the Lorentzian path integral. We have addressed these problems with the following two assumptions: (i) that the energy gap between the false vacuum and the potential top is much smaller than the energy density at the false vacuum and (ii) that the kinetic energy of the homogeneous scalar field is much smaller than the energy density at the false vacuum. These assumptions result in the closed dS background and the simplified equation of motion Eq. (28).
@@
In this paper, we only considered one example with the linear potential. Detailed evaluations, e.g., more complicated potential barrier, for up-tunneling transitions in the Lorentzian path integral need further investigation. We leave this as a future work. Also, throughout this paper, we considered homogeneous tunneling in dS spacetime driven by the quantum tunneling effect, given that the wavefunction is not decohered even at the superhorizon scale. However, it is nontrivial whether the quantum nature of the long-wavelength fluctuations is relevant to the dynamics of dS-horizon patches. This also needs to be investigated by further discussion and computation on the evolution of the primordial fluctuation from other points of view, e.g. quantum information theory.
@@ Acknowledgements
The work of D.S. was supported by JSPS KAKENHI Grants No. 24KJ1223. The work of N.O. was supported by Japan Society for the Promotion of Science KAKENHI Grant No. JP23K13111 and by the Hakubi project at Kyoto University.

### v2

@@ I Introduction
Vacuum phase transition is one of the most important phenomena in cosmology. The decay of a false vacuum is a quantum-tunneling process of a field $\phi$ from a metastable false vacuum to a stable true vacuum in the effective potential $V(\phi)$ at lower or zero temperatures. This type of phase transition was first discussed in Refs. [1, 2], where the transition rate was evaluated by computing the on-shell Euclidean action, which corresponds to the non-trivial configuration of the field in the Wick-rotated space. Most of the subsequent studies have been conventionally following the Euclidean path integral proposed in Refs. [1, 2]. One of the simplest models of the transition in curved spacetime derived using the Euclidean path integral method is the Coleman-De Luccia (CDL) bounce [3], which respects the $O(4)$ symmetry. When a positive large cosmological constant governs the background and a potential barrier is shallower, stochastic fluctuations can be dominant in the transition process. Indeed, in de Sitter (dS) background, the quantum fluctuations of long wavelength modes in a dS patch are squeezed and may be treated as stochastic fluctuations [4, 5, 6, 7], which is often referred to as ”the quantum-to-classical transition” of fluctuation. The most famous instanton describing such a transition is known as the Hawking-Moss (HM) instanton [8, 9]. The transition amplitude of the HM bounce can be indeed reproduced by solving the Fokker-Planck equation in an equilibrium condition [10]. Such a process is often interpreted as a thermal hopping process [11, 12, 13, 14].
@@
In Ref. [11], it was argued that the HM amplitude can be understood as the Boltzmann factor, i.e., a probability of finding the homogeneous configuration at the top of maxima in the associated temperature [11, 12, 13]. This interpretation should be valid only after the whole configuration is in a thermal equilibrium state. For the remote HM transition, it is not clear which thermal or quantum-tunneling effect is more important. As a non-equilibrium state, which could be involved in the case (I) as we will discuss later, is hard to describe in general, we here consider the contribution of the quantum tunneling effect to the remote HM transition. Also, the thermal effect would be suppressed when we address the issue (II). We will then study issues (I) and (II) by considering only the tunneling effect due to the long-wavelength quantum fluctuation while turning off the thermal/stochastic noises.
In this paper, we discuss the two issues raised above by utilizing the Lorentzian path integral with the Picard-Lefschetz theory. Recently, a method to evaluate the transition amplitude in the Lorentzian path integral was proposed in the context of quantum cosmology [18]. In Ref. [18], the authors applied the Picard-Lefschetz theory to find the relevant saddle point in the action, which is important to make the integration convergent. This important technique to perform the Lorentzian path integral was also applied to the radio astronomy [19], the tunneling problem in quantum mechanics [20, 21, 22, 23, 24], the thin-shell vacuum decay process [25] and so on ([26, 27, 28]). We use this technique to see the quantum tunneling effect in the up-tunneling process. One may think that the quantum tunneling effect on the homogeneous transition is negligible as the primordial fluctuation may be decohered on the superhorizon scale. However, the decoherence effect may highly depend on how we split the whole system into an observed system and environment, which involves some artificial choice, more or less. In this sense, the quantum nature of the fluctuation may not vanish even in the superhorizon, and it could affect the subhorizon dynamics. The detailed discussion on this is beyond our scope in this paper.
@@ II Euclidean formalism: CDL bounce and HM bounce
where we have used Eq. (14). Eq. (18) implies that the transition amplitude of the HM bounce is governed by the Boltzmann factor associated with the Bekenstein-Hawking temperature of the dS horizon. From this perspective, the transition amplitude for the HM bounce can be understood as the probability of finding the homogeneous configuration of the scalar field at the maxima of the potential in the thermal system of $T=T_{H}$. In that sense, the HM bounce is often interpreted as a ”thermal” transition. Also, as this Boltzmann factor was obtained via the Fokker-Planck equation with the stochastic noise of a homogeneous field in the dS spacetime [10], we assume that the stochastic noise can be identified as the thermal fluctuations. On the other hand, to get some insight into the remote HM bounce Ref. [11, 12], we will here consider the quantum tunneling effect induced by the quantum fluctuation with superhorizon scale, which is distinct from the thermal (or stochastic) effect, which appears only when interactions between IR and UV modes are turned on [29].
@@ IV.2 Classical action and saddle point
Before presenting our results, it should be noted how we model the remote HM transition in our setup. According to Ref. [11], the HM amplitude can be interpreted as a transition amplitude when we consider the transition from a false vacuum to the neighboring potential top and the potential is sufficiently flat near the top ($\gamma\ll 1$). Additionally, the previous work such as Ref. [10] assumes that the system is in equilibrium, and then interpret the HM transition as a stochastic process. To ensure the system is thermalized within the relevant time scale, say $t$, the field distance between the false vacuum and the potential top $\Delta\phi$ should not be so long that the field cannot reach the top from the false vacuum within $t$. Assuming that the typical relaxation time to achieve equilibrium state, $t_{\rm eq}$, can be estimated by the time scale of the stochastic dissipation process of the field, one could use the following relation [31, 32, 33]
@@
where $H$ is the Hubble parameter and $\sim Ht_{\rm eq}$ is the step number of the stochastic process. One can read $t_{\rm eq}\sim(\Delta\phi)^{2}/H^{3}$ from (38) and find $t_{\rm eq}\propto 1/\gamma$ provided that $\Delta V$ and $H$ are fixed. As the conventional HM transition cannot be applicable to a situation with $t\ll t_{\rm eq}(\Delta\phi,H)$, there should be a certain threshold value $\gamma=\gamma_{\textmd{th}}\ll 1$ above which (i.e., $\gamma_{\textmd{th}}<\gamma\ll 1$) the field space around the false vacuum and potential top achieves equilibrium state within the relevant time scale. On the other hand, with a shallower potential ($\gamma<\gamma_{\textmd{th}}$) and a large $\Delta\phi$, thermal equilibrium is unlikely to be realized around the top of the barrier within the relevant time scale. Thus, if we want to discuss the issue (I) concerning the quantum transition to the remote potential top, the relevant situation is $\gamma<\gamma_{\textmd{th}}$ with $\Delta V$ being fixed  [^1] Recently, the transition rate of the HM transition is reproduced by the stochastic formalism without the assumption of thermal equilibrium [14]. In Ref. [14], the time scale of the transition was evaluated as $t\sim 1/H\gamma$ up to the factor which depends on the detailed shape of the potential. This implies that, for $t\ll t_{\rm eq}$ or $\gamma\ll 1$, even if we relax the condition, the effect of the long wavelength transition can be important.. On the other hand, for issue (II) concerning quantum tunneling with a steeper barrier ($|V^{\prime\prime}|\gtrsim H^{2}$), it is reasonable to consider the region $\gamma\gtrsim 1$, assuming the stochastic noise can be neglected compared to quantum diffusion. In the following, based on the above discussion, we will evaluate the transition amplitudes for the two situations (I) and (II).
@@
Figure 3: The contour of $\text{Re}[iS_{0}(N)]$ and $\text{Im}[iS_{0}(N)]$ for $\Delta V/V_{0}=10^{-6}$. $\gamma=10$. In the left figure, the black markers denote the saddle points, and green/red dotted curves represent the steepest ascent/decent paths, respectively. The black thick line denotes the original contour $(0,\infty)$.
Figure 4: The contour of $\text{Re}[\frac{i}{\hbar}S_{0}(N)]$ with the substitution $\hbar=e^{i\theta}$. The right and left panel denote the cases with $\theta=\frac{\pi}{10}$ and $-\frac{\pi}{10}$, respectively. The value of the parameters are fixed same as in Fig. 3. We can see that the degeneracy of the steepest ascent/decent paths are resolved and the steepest ascent path from the lower saddle point crosses with the original contour $(0,\infty)$.
@@
We present the contour plots of $iS_{0}$ in Fig. 3. We can see that saddle points exist in the complex-$N$ plane. We find two saddle points close to the original contour at $(0,\infty)$ on the real-$N$ axis. Those saddle points are marked in the left top panel of Fig. 3. In the following, we shall denote the saddle points in the upper and lower half region in the complex-$N$ plane as the upper and lower saddle points, respectively. In the left top panel, the steepest ascent (descent) curves from the saddles are plotted as the green (red) dotted lines. We observe that one of the steepest descent curves from the upper saddle point coincides with one of the steepest ascent curves from the lower one. This prevents us from making the one-to-one correspondence between the saddle point and the associated steepest descent and ascent contours, and decomposing the Lefschetz thimbles uniquely. To resolve this problem, the deformation of the path is useful, as discussed in Refs. [18, 34]. Indeed, by taking $\hbar$ to $e^{i\theta}$, as often done in the context resurgence, we can see the degeneracy. In Fig. 4, the value of the real part of $\frac{i}{\hbar}S_{0}$ with $\hbar=e^{i\theta}$ is depicted. We can see that the only one Lefschetz thimble crosses with each saddle point and the steepest ascent curve from the lower saddle point crosses with the original contour $(0,\infty)$, regardless of the signature of $\theta$. Thus, only the lower saddle point contributes to the dominant transition.
@@ IV.4 Transition rate
In both cases, $\Delta V/V_{0}=10^{-6}$ and $10^{-3}$, shown in Fig. 6, the exponent $\text{Im}(S_{\rm eff})$ is equal to or higher than that of the HM bounce for $\gamma<\gamma_{\textmd{cr}}$. As $\gamma\ll 1$ corresponds to the remote HM transition, it is subtle if the conventional HM amplitude $e^{-{\cal B}_{\rm HM}/\hbar}$ gives the transition amplitude in that limit as discussed in Ref. [11]. According to Ref. [11], for the remote HM transition, the conventional HM amplitude merely means the probability of finding the homogeneous configuration, but this says nothing about the relative likelihood with the initial configuration. Thus, when we focus on the homogeneous transition from the false vacuum to the potential top, especially the distant one, the rate may be determined by the quantum tunneling amplitude rather than the Boltzmann factor as the dominant contribution. If this is the case, the remote HM transition would be suppressed as one can see in Fig. 6. For instance, the exponent of the rate is $\text{Im}[S_{\rm eff}]/{\cal B}_{\rm HM}\simeq 89$ for $\gamma=0.1$ and $\Delta V/V_{0}=10^{-6}$. Again, note that the stochastic noise, caused by the interaction between UV and IR modes, is turned off in our computation as we are interested in the short-time transition ($t\ll t_{\rm eq}$) for which the stochastic noise would not play an important role. As such, the deviation from the conventional HM amplitude at $\gamma\sim\gamma_{\rm cr}$ in Fig. 6 does not imply serious discrepancy or conflict.
@@ V Conclusions and discussions
In this paper, we investigated homogeneous tunneling using the Lorentzian path integral, following the Picard-Lefschetz theory, and evaluated the tunneling probability of a homogeneous scalar field. We then discussed the following two issues: (I) the contribution of the quantum tunneling effect to the remote HM transition and (II) the HM transition with a steeper barrier for which the thermal effect would be suppressed. To this end, we turned off thermal/stochastic noises and took into account only the effect of the quantum tunneling with long wavelength by using the Lorentzian path integral. We have addressed these problems with the following two assumptions: (i) that the energy gap between the false vacuum and the potential top is much smaller than the energy density at the false vacuum and (ii) that the kinetic energy of the homogeneous scalar field is much smaller than the energy density at the false vacuum. These assumptions result in the closed dS background and the simplified equation of motion Eq. (28).
@@
In this paper, we only considered one example with the linear potential. Detailed evaluations, e.g., more complicated potential barrier, for up-tunneling transitions in the Lorentzian path integral need further investigation. We leave this as a future work. Also, throughout this paper, we considered homogeneous tunneling in dS spacetime driven by the quantum tunneling effect with long wavelength fluctuation, given that the wavefunction is not decohered even at the superhorizon scale. However, it is nontrivial whether the quantum nature of the long-wavelength fluctuations is relevant to the dynamics of dS-horizon patches. This also needs to be investigated by further discussion and computation on the evolution of the primordial fluctuation from other points of view, e.g. quantum information theory.
@@ Acknowledgements
D.S. is grateful to T. Miyachi, K. Okabayashi, and J. Tokuda for useful comments and fruitful discussion. The work of D.S. was supported by JSPS KAKENHI Grants No. 24KJ1223. The work of N.O. was supported by Japan Society for the Promotion of Science KAKENHI Grant No. JP23K13111 and by the Hakubi project at Kyoto University.